OpenStudy (anonymous):

if the equation of a parabola is y=7(x-4)^2+v, then what is the value of v?

OpenStudy (anonymous):

could be anything.

OpenStudy (anonymous):

if the vertex is (4, -2) forgot to add that part

OpenStudy (anonymous):

okay, substitute y = -2 and x = 4 in the equation and find v

OpenStudy (anonymous):

I'm kind of confused

OpenStudy (anonymous):

what are you confused about? have you ever done substitution?

OpenStudy (anonymous):

I haven't ever had anyone explain it to me in detail. I'm home schooled and my mom doesn't have any knowledge of this stuff

OpenStudy (anonymous):

okay the equation of the parabola is given. the vertex is given. the vertex is a point on the parabola right? so we can substitute the vertex instead of x and y in the equation so y = 7(x-4)^2+v becomes -2 = 7(4-4)^2+v 4-4 =0, so -2 = 7*0^2 +v -2 = v or v = -2